# Coherence restoring in communication line via controlled interaction with environment

### v1

@@ Abstract
We consider the state-restoring protocol based on the controlled interaction of a linear chain with environment through the specially adjusted step-wise time dependent Lindblad operators. We show that the best restoring result (maximal scale factors in the restored state) corresponds to the symmetrical Lindblad equation. (0,1)-excitation dynamics is considered numerically, and restoring protocol for the 1-order coherence matrix is proposed for the case of the two-qubit sender (receiver). The state-restoring with equal scale factors is also considered reflecting the uniform scaling of the restored information.
Coherence restoring in communication line via controlled interaction with environment
E.B. Fel’dman1, I.D. Lazarev1,2, A.N. Pechen3,4 and A.I. Zenchuk1
1 Federal Research Center of Problems of Chemical Physics and Medicinal Chemistry RAS, Chernogolovka, Moscow reg., 142432, Russia.
2 Faculty of Fundamental Physical-Chemical Engineering, Lomonosov Moscow State University, GSP-1, Moscow, 119991, Russia.
3 Department of Mathematical Methods for Quantum Technologies, Steklov Mathematical Institute of Russian Academy of Sciences, Gubkina str. 8, Moscow 119991, Russia
4National University of Science and Technology MISIS, Leninski prosp. 4, Moscow 119991, Russia
Corresponding author. E-mail: zenchuk@itp.ac.ru
@@
1 Introduction
Contemporary quantum information technology  [1, 2] motivates the intensive development of quantum transmission protocols, which are based on photons [3, 4, 5] or, alternatively, on spin chains [6, 7] when dealing with solid state quantum architectures. The teleportation of arbitrary state [8, 9, 10], arbitrary state transfer [11, 12, 13, 14, 15, 16, 17, 18], remote state creation [3, 4, 5, 6, 7, 19, 20] represent different, but intertwined directions in development of this topic. The intensive development of quantum technologies [21] stimulates further elaboration of optimization technique for quantum information transfer along communication lines. Among numerous results devoted to quantum communication protocols we mention only several of them, such as quantum state transfer through one dimensional rings of qubits with fixed interactions [22], fast high-fidelity information transfer using either a single one-side control [23] or controls involving both ends of the lattice [24], communication at the quantum speed limit [25, 26, 27], many-body state generation [28], functional donor chain [29], information transfer along the chaotically kicked spin chains [30], controlled state transfer along the Heisenberg XXZ spin chains governed via periodic drives [31], local control in non-adiabatic cutting and stitching of a spin chain [32], excitation propagation along a chain with optimized site- dependent interaction strengths [33], universal control and error correction applied to multi-qubit spin registers in nitrogen-vacancy center [34], dipolar spin chains with Floquet prethermalization  [35], coherent control of a nuclear spin dynamics via interactions with a rare-Earth ion [36], etc. Many other aspects concerning controlling protocols in application to spin systems are explored in book [37].
@@ 1 Introduction
In our paper, we further develop the problem of remote restoring the state transferred from the sender $S$ to the receiver $R$ along a spin-1/2 chainRefs. [38, 39, 40, 41, 42, 43, 44]. The nodes of a spin chain connecting sender and receiver form the transmission line $TL$. Initially, the sender is prepared in the(arbitrary) state $\rho^{(S)}(0)$, while all the spins of $TL$ and $R$ are in the ground states. Then the prepared initial state is subjected to the time-evolution governed by some Hamiltonian, effected by the interaction with environment, up to some time instant $t_{\mathrm{reg}}$ which we call time instant for state registration. The initial sender state is restored if, at that time instant, the elements of the receiver’s density matrix $\rho^{(R)}(t_{\mathrm{reg}})$ are proportional to the appropriate elements of the sender’s initial density matrix $\rho^{(S)}(0)$:
@@
The paper is organized as follows. In Sec. 2, we consider the general state-restoring protocol based on the time-depending Lindbladian. Evolution under the time-dependent Lindbladian including XXZ-Hamiltonian is discussed in Sec. 3. State restoring of (0,1)-excitation states with numerical simulation of 8 and 10 node chains and two-qubit sender (receiver) is given in Sec. 4. Concluding remarks are presented in Sec. 5.
2 General state restoring protocol
We consider the evolution of a quantum system interacting with the environment described by the Lindblad equation
@@ 2 General state restoring protocol
where both the Hamiltonian $H$ and the Lindblad operators $L_{i}$ conserve the excitation number in the system. We emphasize that the Lindblad operators are time-dependent which is necessary to establish the control over the restoring process. The state restoring protocol is similar to that proposed in [43], but there is significant difference because of the non-Hamiltonian evolution.
@@
where $|n|$ is the number of units in the binary representation of the index $n$, $U(t)$ is the evolution operator defined by the Lindblad equation (2) with the initial condition
@@
3 State restoring via Lindbladian with XXZ-Hamiltonian
We consider the evolution under the Lindbladian with XXZ-Hamiltonian and $L_{i}(t)=\sqrt{\gamma_{i}(t)}Z_{i}$. Since $Z_{i}^{2}=\frac{E}{4}$ ($E$ is the identity operator), Eq.(2) gets the form
@@ 3 State restoring via Lindbladian with XXZ-Hamiltonian
Here $D_{ij}=\gamma^{2}/r_{ij}^{3}$ are the coupling constants between the $i$th and $j$th spins (for $\hbar=1$), $\gamma$ is the gyromagnetic ratio, $r_{ij}$ is the distance between the $i$th and $j$th spins, and $X_{i},\;\;Y_{i}$ are the operators of, respectively, the $x$- and $y$-projections of the $i$th spin, the external magnetic field is along the spin chain. Of course, commutation condition (8) is satisfied. The block-structure of $L_{i}$ in (10) is related to the block structure of $Z_{i}$:
@@
4 $(0,1)$-excitation states. Restoring 1-order coherence matrix
Now we consider the (0,1)-excitation space and construct the protocol for restoring the elements of the 1-order coherence matrix $\rho^{(1)}$ only, thus leaving all elements of the 0-order coherence matrix unrestored. In this case Eq.(18) gets the following form:
@@
Eq.(3) gets the following form:
@@
($\lambda=1$ in the perfect case). We denote by $\gamma^{\mathrm{(opt)}}$ the list of parameters $\tilde{\gamma}^{(m)}$ found as the result of maximization in (48).
4.1 Numerical simulations, $N^{(S)}=2$
@@
$R_{\mathrm{\lambda}}(\gamma)=|1-\lambda_{01}(\gamma)|^{2}+|1-\lambda_{10}(\gamma)|^{2}.$ (51)
Thus, we solve the system (49) with the regularization functional (51) using the least-squares method implemented in the least_squared function from the SciPy package [65]. The sum of squared residuals has the form
$S(\gamma)=\delta_{1}^{2}(\gamma)+\delta_{2}^{2}(\gamma)+\mu\left(\left(|1-\lambda_{01}(\gamma)|\right)^{2}+\left(|1-\lambda_{10}(\gamma)|\right)^{2}\right)$ (52)
where $\mu$ is a parameter that controls the importance of regularization. In all calculations presented below, we used $\mu=10^{-6}$. Finally, we use the approximate solution obtained at this stage as an initial approximation to find the solution of the restoring system (49) without regularization functional (51), i.e., setting $\mu=0$ in (52).
@@
4.1.1 Centrally-symmetric distribution of damping rates
@@
4.1.2 Edges and center model
@@
4.1.3 State restoring with equal $\lambda$-parameters
@@ 5 Conclusions
The state restoring protocol proposed in this paper uses a special time-dependent interaction with the environment thus replacing the restoring unitary transformation of the extended receiver that was originally used in state-restoring protocols [38, 39, 40, 41, 42]. Although the $\lambda$-parameters are not large by absolute value for this type of restoring, the environment control might be combined with, for instance, the control by the local magnetic field [43] and thus expand the freedom in choice of control parameters.
@@
Acknowledgments. The work was carried out as a part of a state task, State Registration No. 124013000760-0.

### v2

@@ Abstract
We consider the state-restoring protocol based on the controlled interaction of a linear chain with environment through incoherent control by the specially adjusted step-wise time dependent Lindblad operators. We show that the best restoring result (maximal scale factors in the restored state) corresponds to the symmetrical Lindblad equation. (0,1)-excitation dynamics is considered numerically, and restoring protocol for the 1-order coherence matrix is proposed for the case of the two-qubit sender (receiver). The state-restoring with equal scale factors is also considered reflecting the uniform scaling of the restored information.
@@
I Introduction
Contemporary quantum information technology  [1, 2] motivates the intensive development of quantum transmission protocols, which are based on photons [3, 4, 5] or, alternatively, on spin chains [6, 7] when dealing with solid state quantum architectures. The teleportation of arbitrary state [8, 9, 10], arbitrary state transfer [11, 12, 13, 14, 15, 16, 17, 18], remote state creation [3, 4, 5, 6, 7, 19, 20] represent different, but intertwined directions in development of this topic. The intensive development of quantum technologies [21] stimulates further elaboration of optimization technique for quantum information transfer along communication lines.
I.1 Brief review on quantum communication protocols
Among numerous results devoted to quantum communication protocols we mention only several of them. Quantum state transfer through one dimensional rings of qubits with fixed nearest neighbor interactions constant throughout the ring is considered in Ref.[22]. Two parties are positioned at the opposite sites of the ring and both can operate with certain number of their nearest qubits to increase the fidelity of state transfer. Fast high-fidelity information transfer using the one-side control via on-off switching actuator is considered in Ref.[23]. The effective switching sequence can increase both speed and fidelity of state transfer, the particular form of the Hamiltonian has no principal role. The two side control of high-fidelity state transfer along the disordered spin chain governed by the nearest-neighbor Hamiltonian including only $x$-directed spin-spin interaction is proposed in Ref.[24]. The state-transfer optimization tool in this case includes two Larmor frequencies at the end spins of the chain. The control of state transfer by the pair of time-dependent end-bonds in the XX-Hamiltonian with nearest-neighbor interaction is considered in Ref.[25]. The two-order magnitude increase in the rate of state transfer governed by Heisenberg-type Hamiltonian can be reached via the optimal control of the external parabolic magnetic field reaching the theoretical maximum of propagation rate which is $\sim N/\langle D\rangle$ ($\langle D\rangle$ is the average coupling constant in the Hamiltonian), as shown in Ref.[26]. Speed limits for evaluation of one- two- and three- qubit operations and state transfer depending on the type of spin-interaction and geometry is studied in Ref.[27]. The protocol for preparing the multiqubit spin chain in the needed state using the optimized coherent pulses, engineered dissipation and feedback is proposed in Ref.[28]. The state transfer along the odd-sized donor chain with realistic conditions is studied in Ref.[29], where the nearest-neighbor coupling constants and noise level are optimized for the high-fidelity state transfer. The spin dynamics and information transfer among the spins of the chain governed by either Heisenberg or Ising-Z -Hamiltonian where each spin is subjected to a chaotic kicks is studied in Ref.[30]. State transfer along the chain governed by the Heisenberg XXZ-Hamiltonian with periodic driving potential leads to either perfect state transfer or to complete blockade of the input signal, Ref.[31]. This feature can be used in formation of quantum logic elements. The local control based on non-adiabatic cutting and stitching of end-bonds in a spin chain using special time-dependent driving function can be also implemented in quantum information technology, Ref.[32]. In Ref.[33], it is demonstrated that almost perfect single excitation state transfer is possible using a spin chain governed by the Heisenberg Hamiltonian with either nearest-neighbor interactions, or involving remote spin interactions with optimized site-dependent coupling constants. Non-trivial Floquet phases of matter are obtained in the periodically driven Floquet quantum chain with special engineering, Ref.[34]. The control of a nuclear spin coherently coupled with the electron spins of rare-Earth ions is studied in Ref. [35] in application to quantum repeater. In this case the control is established by the special sequence of magnetic pulses applied to the electron spin. The control via weakly coupled nuclear spins and electron spins in the nitrogen-vacancy center in diamond is applied to multiqubit spin registers in Ref.[36] allowing to prepare the high-fidelity single- and two-qubit gates with error correction protocol. Many other aspects concerning controlling protocols in application to spin systems are explored in book [37].
I.2 Controlled interaction with environment
In this work, we exploit controlled interaction with the environment for coherence restoring in communication qubit chain lines. While dissipation and interaction with the environment are often considered as having deleterious effect on the abilities to manipulate quantum systems, in some cases they can be exploited as a useful resource for controlling quantum systems as was shown in the approach of [38], where a general incoherent control method for manipulation by quantum systems exploiting controlled interaction of the system with generally time-dependent environment and induced by this interaction time-dependent master equations was developed for various physical forms of the environment. In this approach, a quantum system interacts with a time-dependent environment and this interaction leads to a master equation with time-dependent coefficients and with engineered dissipation. The master equation in particular can have the Gorini–Kossakowski–Sudarshan–Lindblad (GKSL) form but with time-dependent coefficients; however, non-Markovian master equations can be considered as well. Important that the time-dependent master equations in [38] were exploited not in the abstract form but in various explicit forms corresponding to various physical environments. Two general types of the environment were considered as explicit examples—environment former by incoherent photons or phonons (Sec. II in [38]) and collision type environments (e.g., gas, spin particles, etc.) which interact with the controlled system via collisions (Sec. III in [38]). The first type of environment corresponds to the weak coupling and rotating wave Markovian approximation and the second type of the environment corresponds to the low density Markovian approximation. For both cases, explicit form of time-dependent Lindblad operators was written and exploited for quantum control. In this way, generally time-dependent state of the environment is used as a control to adjust the time-dependent decoherence rates and control evolution of the system density matrix. This incoherent control approach was shown to be quite powerful to provide, when combined with coherent control by lasers or electromagnetic field, creation of arbitrary mixed density matrices of generic quantum systems [39].
Controlled interaction with the environment and engineered dissipation were used for engineering arbitrary pure and mixed quantum states of generic quantum systems [39] which are necessary in some models of quantum computation with mixed states and non-unitary gates [40, 41, 42, 43], in nuclear magnetic resonance [44], steady quantum states generation [45], for improving quantum state engineering and computation [42], cooling of translational motion without reliance on internal degrees of freedom [46], preparing entangled states [47, 48], inducing multi-particle entanglement dynamics [49], making robust quantum memories [50], creating entanglement dynamics in two-qubit systems [51], and in other applications. Engineered interaction with the environment was used for making pointer-like states by suitably engineering the coupling to the environment via open-loop periodic control [52], dissipatively preparing arbitrary pure quantum states on a multipartite qubit register in a finite number of basic control blocks [53], deterministically generating entanglement in an ensemble of strongly interacting Rydberg atoms [54], preparing a spin chain in a generic many-body state in the asymptotic limit of tailored nonunitary dynamics [28], optimization of up-conversion hues in phosphors [55], preparation of qubit superposition states as the steady state of coherent driving and dissipation near the photonic crystal band edge [56], manipulation by two-spin systems [57], driven-dissipatively preparing and stabilizing a quantum manybody entangled state with symmetry-protected topological order [58], generation of target states and quantum gates in open quantum systems [59, 60]. Many more examples can be found in review [61].
I.3 State-restoring algorithms
@@
We study the restoring of the state transferred from the sender $S$ to the receiver $R$ along a spin-1/2 chain [62, 63, 64, 65, 66, 67, 68]. The nodes of a spin chain connecting sender and receiver form the transmission line $TL$. Initially, the sender is prepared in the (arbitrary) state $\rho^{(S)}(0)$, while all the spins of $TL$ and $R$ are in the ground states. Then the prepared initial state is subjected to the time-evolution governed by some Hamiltonian, effected by the interaction with environment, up to some time instant $t_{\mathrm{reg}}$ which we call time instant for state registration. The initial sender state is restored if, at that time instant, the elements of the receiver’s density matrix $\rho^{(R)}(t_{\mathrm{reg}})$ are proportional to the appropriate elements of the sender’s initial density matrix $\rho^{(S)}(0)$:
@@ I.4 Optimization methods
The paper is organized as follows. In Sec. II, we consider the general state-restoring protocol based on the time-depending Lindbladian. Evolution under the time-dependent Lindbladian including XXZ-Hamiltonian is discussed in Sec. III. State restoring of (0,1)-excitation states with numerical simulation of 8 and 10 node chains and two-qubit sender (receiver) is given in Sec. IV. Concluding remarks are presented in Sec. V.
II General state restoring protocol
We consider the Born-Markov approximation for evolution of a quantum system interacting with the environment is described by the master equation of the general GKSL form
@@ II General state restoring protocol
A rigorous derivation of the general form of Markovian master equations was performed in Refs.[85, 86, 87]. For a simple physical justification one can look, for example in [88]. Time-dependent form of the master equation with explicit forms of the time-dependent Lindblad operators $L_{i}(t)$ corresponding to two physical models of the environment, namely environment formed by photons or phonons (corresponding to weak coupling and rotating wave markovian approximation), and environment formed by collisional type media (quantum gas, spin systems, etc., corresponding to low density approximation) was considered in [38].
We consider an additional requirement that both the Hamiltonian $H$ and the GKSL operators $L_{i}$ conserve the excitation number in the system. A particular realization of such Lindblad operators is proposed below in Sec. III. The requirement on conserving the excitation number is auxiliary, it is included to simplify description of the spin dynamics. In our case, this dynamics will be described by two block of the Lindblad operator responsible for 0- and 1-excitation evolution. In general, Eq.(II) holds for any Hamiltonian and Lindblad operators. We emphasize that the Lindblad operators are time-dependent which is necessary to establish the control over the restoring process. The state restoring protocol is similar to that proposed in [67], but there is significant difference because of the non-Hamiltonian evolution.
@@
where $|n|$ is the number of units in the binary representation of the index $n$, $U(t)$ is the evolution operator defined by the Lindblad equation (II) with the initial condition
@@
III State restoring via Lindbladian with XXZ-Hamiltonian
We consider the evolution under the Lindbladian with XXZ-Hamiltonian and such interaction with the environment that results only in dephasing in the spin system. This type of interaction can be described by the Lindblad operators $L_{i}(t)=\sqrt{\gamma_{i}(t)}Z_{i}$. Since $Z_{i}^{2}=\frac{E}{4}$ ($E$ is the identity operator), Eq.(II) gets the form
@@ III State restoring via Lindbladian with XXZ-Hamiltonian
Here $D_{ij}=\gamma^{2}/(2r_{ij}^{3})$ are the coupling constants between the $i$th and $j$th spins (for $\hbar=1$), $\gamma$ is the gyromagnetic ratio, $r_{ij}$ is the distance between the $i$th and $j$th spins, and $X_{i},\;\;Y_{i}$ are the operators of, respectively, the $x$- and $y$-projections of the $i$th spin, the external magnetic field is perpendicular to the spin chain. Of course, commutation condition (8) is satisfied. We note that the rotating-wave approximation [89], which was used in deriving the XXZ-Hamiltonian, is represented by the unitary transformation $e^{i\gamma BI_{z}}$ ($B$ is the intensity of the external magnetic filed) commuting with our Lindblad operators and therefore does not change them.
The block-structure of $L_{i}$ in (10) is related to the block structure of $Z_{i}$:
@@
IV Restoring 1-order coherence matrix
Now we consider the (0,1)-excitation state-space, i.e., the state-space spanned by the eigenvectors with 0- and 1 excitations, and construct the protocol for restoring the elements of the 1-order coherence matrix $\rho^{(1)}$ only, thus leaving all elements of the 0-order coherence matrix unrestored. In this case Eq.(18) gets the following form:
@@ IV Restoring 1-order coherence matrix
Eq.(III) gets the following form:
@@
($\lambda=1$ in the perfect case). We denote by $\gamma^{\mathrm{(opt)}}$ the list of parameters $\tilde{\gamma}^{(m)}$ found as the result of maximization in (48). We emphasize that the parameter $\lambda$ in Eq.(48) indicates how much the absolute values of the matrix elements of $\rho^{(S)}(0)$ are decreased during the state transfer and restoring. This is the most important characteristics of restoring process because vanishing $\lambda$ means loosing the transferred information. However, the phases of the $\lambda$-parameters are also important because they are added to the phases of the elements of $\rho^{(S)}(0)$. Therefore, both absolute values and phases of the $\lambda$-parameters must be kept in the state-restoring protocol, see Refs.[62, 63, 64, 65, 66].
IV.1 Numerical simulations, $N^{(S)}=2$
@@
$R_{\mathrm{\lambda}}(\gamma)$ $=$ $\left(1-|{\mbox{Re}}(\lambda_{01}(\gamma))|\right)^{2}+\left(1-|{\mbox{Im}}(\lambda_{01}(\gamma))|\right)^{2}+$
$\left(1-|{\mbox{Re}}(\lambda_{10}(\gamma))|\right)^{2}+\left(1-|{\mbox{Im}}(\lambda_{10}(\gamma))|\right)^{2}.$
Thus, we solve the system (49) with the regularization functional (IV.1) using the least-squares method implemented in the least_squared function from the SciPy package [90]. The sum of squared residuals $\delta_{1}^{2}$ and $\delta_{2}^{2}$ and weighted regularization function has the form
$S(\gamma)=\delta_{1}^{2}(\gamma)+\delta_{2}^{2}(\gamma)+\mu{R_{\mathrm{\lambda}}(\gamma)}$ (52)
where $\mu$ is a parameter that controls the importance of regularization. In all calculations presented below, we used $\mu=10^{-6}$. Finally, we use the approximate solution obtained at this stage as an initial approximation to find the solution of the restoring system (49) without regularization functional (IV.1), i.e., setting $\mu=0$ in (52).
@@
IV.1.1 Centrally-symmetric distribution of damping rates
@@
IV.1.2 Edges and center model
@@
IV.1.3 State restoring with equal $\lambda$-parameters
@@ V Conclusions
The state restoring protocol proposed in this paper uses a special time-dependent interaction with the environment thus replacing the restoring unitary transformation of the extended receiver that was originally used in state-restoring protocols [62, 63, 64, 65, 66]. Although the $\lambda$-parameters are not large by absolute value for this type of restoring, the environment control might be combined with, for instance, the control by the local magnetic field [67] and thus expand the freedom in choice of control parameters. In addition, the absolute values of the $\lambda$-parameters can be increased using any special tool increasing the fidelity of state transfer, for instance, the week end-bonds [14, 91, 64]. However, even the absolute value $\sim 0.1$ obtainable in this paper can be acceptable, for instance, in the measurement based algorithm for the perfect state transfer [92].
@@
Acknowledgments. The work was carried out as a part of a state task, State Registration No. 124013000760-0. A.P. acknowledges partial support of the state task at MIAN.